Find the equation of a line having the given slope and containing the given point.
m=8/9, (2,-6)

Answer by ewatrrr(24785)

 
Hi,


m = 8/9
y = (8/9)x + b |using ordered pair(2,-6) to solve for b
-6 = (8/9)*2 + b
-6 -16/9 = b
-70/9 = b
y = (8/9)x - 70/9